# Compelling ReLU Networks to Exhibit Exponentially Many Linear Regions at Initialization and During Training

### v4

@@ Appendix A Appendix
Figure 9: Extending our results into a larger network. ReLU should not be used in the layers between the blocks, as creating nonconvex functions this way will tend to make a lot of negative values.
@@ A.1 Initialization and Training Algorithm
The network weights can then be initialized once more based on the learned vector $A$ so that the network weights can be updated via regular gradient descent. In our experiments, both phases of training (pretraining and gradient descent) ran for 1000 epochs. Full pseudocode is listed in Algorithm 1.
A.2 Derivations
For convenience we first restate the functions defined in the main body of the paper.
@@ A.2 Derivations
Notationally, we will denote the sorted $x$-locations of the peaks and valleys of $W_{i}(x)$ by the lists $P_{i}=\{x:W_{i}(x)=1\}$ and $V_{i}=\{x:W_{i}(x)=0\}$. We will use the list $B_{i}$ to reference the locations of all non-differentiable points, which we refer to as bends. $B_{i}:=P_{i}\cup V_{i}$. $f_{i}(x)=\sum_{n=0}^{i-1}s_{n}W_{n}(x)$ will denote finite depth approximations up to but not including layer $i$. The error function $E_{i}(x)=\sum_{n=i}^{\infty}s_{n}W_{n}(x)=F(x)-f_{i}(x)$ will represent the error between the finite approximation and the infinite depth network. Although the derivative of an infinite sum of functions is not always equal to the sum of the individual functions derivatives, it is in this case, so notationally we use $F^{\prime}$ to describe the latter for brevity.
@@
The goal of this section will be to show how to select the $s_{i}$ based on $a_{i}$ in a manner where the derivative $F^{\prime}(x)$ is defined and on all of $[0,1]$. Fortunately, the left and right derivative limits $F^{\prime}_{+}(x)$ and $F^{\prime}_{-}(x)$ already exist, since each bend in each $W_{i}$ will use the slope of the line segment to its left or right respectively. The problem is that without choosing the scaling parameters appropriately, $F^{\prime}_{+}(x)$ and $F^{\prime}_{-}(x)$ will take different values so that $F^{\prime}(x)$ will be undefined on all bend points.
@@
Here we compute the right derivative $(E_{i})^{\prime}_{+}(x)$ of the error function at a point $x$. The left derivative will only be different by a constant factor.
@@
Let $x_{n}$ be some point in $P_{i}$, and let $n$ be its index in any list it appears in. To calculate the value of $E_{+}^{\prime}(x)=\sum_{n=i}^{\infty}s_{n}(W_{n})^{\prime}_{+}(x)$, we will have to find the slope of the linear intervals to the immediate right of $x_{n}$ for all $W_{i}$. The terms in the summation Equation 5 mostly derive from the chain rule. We will use $R$ to represent $W_{i+}^{\prime}(x)$. The first term in the sum will be $Rs_{i}$. $W^{\prime}_{i+1}(x)=T^{\prime}_{i+1}(W_{i}(x))W^{\prime}_{i}(x)$ since the derivatives of composed functions will multiply. There are two different possible slope values of $T_{i+1}$ to multiply by, and the correct one to multiply is $-1/(1-a_{i+1})$ because $x_{n}$ is a peak of $W_{i}$, so $W_{i}(x)>a_{i+1}$ for $x\in(B_{i+1}[n-1],B_{i+1}[n+1])$. This gives $W^{\prime}_{i+1}(x)=-R\frac{s_{i+1}}{(1-a_{i+1})}$. Note that the second term has the opposite sign as the first.
For all remaining terms, since $x_{n}$ was in $P_{i}$, it is in $V_{j}$ for $j>i$. For $x\in(B_{j+1}[n-1],B_{j+1}[n+1])$, $W_{j}(x)<a_{j+1}$ and the chain rule applies the first slope $1/a_{j+1}$. Since this slope is positive, every term has the opposite sign as the first. Summing up all the terms with the coefficients $s_{i}$, and factoring out $R$ will yield the desired formula. Note that this same reasoning applies to the left side, the initial slope $L$ will just be different. ∎
@@
This lemma shows that to calculate the derivative at of $F(x)$ for any bend point $x$ (whenever it’s actually defined), one needs only to compute the derivative of the finite approximation $f_{i}$ (which excludes $W_{i}$). This will be useful later for proving other results. A visualization of this lemma is provided in Figure 11
@@
From the previous lemma we have $E^{\prime}(x)=0$ whenever the directional derivatives are equal. $F(x)=\sum_{j=0}^{i-1}s_{j}W_{j}(x)+E(x)$. The first $i-1$ terms are differentiable at the points $P_{i}$ since those points lie between the discontinuities in $B_{i-1}$. Therefore $F^{\prime}(x)$ is defined and can be calculated using the finite sum. ∎
We now prove our main theorem. Intuitively, the theorem shows that there is a way to sum the triangular waveforms $W_{i}$ so that the resulting function approximation converges to a differentiable function, which, as mentioned at the start of this subsection, can aid in preventing overfitting when using exponentially many linear regions. The idea of the proof is that much of the formula for $E^{\prime}(x)$ will be shared between two successive generations of peaks. Once they are both valleys, they will behave the same, so the sizes of their remaining discontinuities will need to be proportional.
@@
$F^{\prime}(x)$ is defined on $[0,1]$ only if the scaling coefficients are selected based on $a_{i}$ according to:
@@
A.3 Sufficiency for differentiability
We can show that in addition to being necessary for the existence of the derivative on the bend points, our choice of scaling is sufficient when $a_{i}$ are bounded away from 0 or 1.
@@ A.3 Sufficiency for differentiability
$F^{\prime}(x)$ is well-defined if we can find $c>0$ such that $c\leq a_{i}\leq 1-c$ for all $i$ and choose all $s_{i}$ according to Equation 4.
@@
We will prove our result by substituting Equation 4 into this formula, and then verifying that the resulting equation is valid. First we would like to rewrite each $s_{n}$ in terms of $s_{i}$. Equation 4 gives a recurrence relation. Converting it to an non-recursive representation we have:
@@
This equation has an interpretation that is important to the argument. 1 is the full size of the initial discontinuity at a point in $P_{i}$, and each term on the other side represents proportionally how much the discontinuity is closed for each triangle wave that is added. Every time a wave is added, it can be thought of as subtracting from each side the first term appearing on the right. We will now show that each term of the sum on the right accounts for a fraction (equal to $a_{i}$) of the remaining error. Inductively we can show:
@@
This means that as the first term appearing on the right is repeatedly subtracted, that term is always equal to $a_{n}$ times the left side. As a base case, we have $(1-a_{i+2})=(1-a_{i+2})$. Assuming the above equation holds for all previous values of $j$
@@ A.4 Error Decay
A.5 Second Derivatives
Here we show that any function represented by one of these networks that is not $y=x^{2}$ does not have a second derivative defined at the bend locations. To show this we will sample a discrete series of $\Delta y/\Delta x$ values from $F^{\prime}(x)$ and show that the limits of these series on the right and left are not the same (unless all $a_{i}=0.5$), which implies that $F^{\prime\prime}(x)$ does not exist (see Figure 11 below). First we will produce the series of $\Delta x$. Let $x$ be the location of a peak of $W_{i}$, and let $l_{n}$ and $r_{n}$ be its immediate neighbors in $B_{i+n}$.
@@ A.5 Second Derivatives
Figure 11: The derivatives of the first few stages of approximation. Notice that each time a constant segment“splits”, the two neighboring segments adjacent to the split converge back to the original value. The points in the series approaching $x$ from the right are marked. $a_{0}\neq 0.5$ and all other parameters are set to $0.5$, which will cause the left and right sets of points to lie on lines with different slopes.
@@
The points $l_{n}$ and $r_{n}$ are all peak locations, Equation 6 gives their derivative values as $f^{\prime}_{i+n}(r_{n})$. In our argument for sufficiency, we reasoned about the sizes of the discontinuities in $f^{\prime}$ at $x$. Since $l_{n}$ and $r_{n}$ always lie on the linear intervals surrounding $x$ as $n\rightarrow\infty$, we can get the value of $f^{\prime}_{i}(x)-f^{\prime}_{i+n}(r_{n})$ using Equation 9 with the initial discontinuity size set to $R*s_{i}$ rather than 1. Focusing on the right hand side we get:
@@
The issue which arises is that the derivation on the left is identical, except for a replacement of $R$ by $L$. The only way for these formulas to agree then is for $R=L$ which implies $a_{i}=1-a_{i}=0.5$. Since this argument applies at any layer, then all $a$ parameters must be $0.5$ (which approximates $y=x^{2}$. ∎
A.6 Concavity
Each of the $f^{\prime}_{i}$ are composed of constant value segments, we will show that those values are monotonically decreasing (this can be seen in Figure11). This then allows us to show that $F$ is concave.
@@ A.6 Concavity
We establish two inductive invariants. One is that the y-values in the list $Y_{i}$ remain sorted in descending order. The other is that $Y^{\prime}_{i}[n]\geq f^{\prime}_{i}(x)\geq Y^{\prime}_{i}[n+1]$ for $V_{i}[n]<x<V_{i}[n+1]$, indicating that the constant value segments of $f_{i}$ lie in between the limits in the list $Y_{i}$. Together, these two facts imply that each iteration of the approximation $f_{i}$ is concave. Which then can prove that their limit $F$ is also concave.
@@
A.7 Learning Rates

### v5

@@ Appendix A Appendix
Figure 9: Extending our results into a larger network. ReLU should not be used in the layers between the blocks. Colors correspond to Figure 2.
@@ A.1 Initialization and Training Algorithm
The network weights can then be initialized once more based on the learned vector $A$, to then update the weights via regular gradient descent. In our experiments, both phases of training (pretraining and gradient descent) ran for 1000 epochs. Full pseudocode is listed in Algorithm 1.
A.2 Terminology
Before presenting further formal mathematical details of our method, we first briefly review a few pieces of basic ReLU network terminology used in the paper. The reader is referred to the Chmielewski-Anders (2020) for an excellent alternative presentation of these terms.
An activation pattern is a boolean mask that tracks which neurons in a network have their output zeroed by ReLU activations. The activation regions of a ReLU network are connected (and can be shown to be convex) sets of inputs on which the activation pattern is constant. Since the action of the ReLU activation function is constant, the network output over an activation region is equivalent to the case where there is no activation function and the associated zeroed neurons are absent; therefore, the network output behaves linearly over the inputs in the activation region. Relatedly, a linear region is a set of inputs on which the network output behaves linearly with respect to its inputs. It may consist of multiple neighboring activation regions. Another important concept is the boundary of a neuron, as described in Rolnick & Kording (2020): the set of inputs for which the neuron outputs 0, independently of ReLU. The boundary of a neuron is precisely the boundary of the activation regions it adds to the network. Chen & Ge (2024) refers to this set as the characteristic activation boundary since these are the boundaries of the activation regions.
We note that although works like ours are theoretical papers that leverage these concepts, studying these ideas can lead to interesting applied learning research. For instance, Rolnick & Kording (2020) leverages theoretical works like those cited in our related work discussion for an exciting privacy application: it is proven that a ReLU network’s output can often be provably used to reverse engineer the architecture of a ReLU network, up to isomorphism.
A.3 Mathematical Results
For convenience, we first restate the functions defined in the main body of the paper.
@@ A.3 Mathematical Results
The goal of this section will be to show how to select the $s_{i}$ based on $a_{i}$ in a manner where the derivative $F^{\prime}(x)$ is defined and continuous on all of $[0,1]$. We begin by assuming that
$F^{\prime}(x)=\sum_{i=0}^{\infty}s_{i}W^{\prime}_{i}(x)$
We will see that the resulting choice of $s_{i}$ ensures uniform convergence of the derivative terms, so that the derivative of the infinite sum is indeed infinite sum of the derivatives. Fortunately, the left and right derivatives $F^{\prime}_{+}(x)$ and $F^{\prime}_{-}(x)$ already exist everywhere, since each bend in each $W_{i}$ (where the full derivative of $F$ is undefined) is assigned the slope of the line segment to its left or right respectively. The $s_{i}$ scaling values will have to be chosen appropriately so that $F^{\prime}_{+}(x)$ and $F^{\prime}_{-}(x)$ are equal for all bend points.
Notationally, we will denote the sorted $x$-locations of the peaks and valleys of $W_{i}(x)$ by the lists $P_{i}=\{x:W_{i}(x)=1\}$ and $V_{i}=\{x:W_{i}(x)=0\}$. We will use the list $B_{i}$ to reference the locations of all non-differentiable points, which we refer to as bends. $B_{i}:=P_{i}\cup V_{i}$. $f_{i}(x)=\sum_{n=0}^{i-1}s_{n}W_{n}(x)$ will denote finite depth approximations up to but not including layer $i$. The error function $E_{i}(x)=\sum_{n=i}^{\infty}s_{n}W_{n}(x)=F(x)-f_{i}(x)$ will represent the error between the finite approximation and the infinite depth network. This odd split around layer $i$ makes the proofs cleaner.
@@
Thanks to the previous lemma, we only need to work with $E_{i}$. Here we compute the right derivative $(E_{i})^{\prime}_{+}(x)$ of the error function at a point $x$. The left derivative will only be different by a constant factor.
@@
Let $x_{k}$ be some point in $P_{i}$, and let $k$ be its index in any list it appears in. To calculate the value of $E_{+}^{\prime}(x_{k})=\sum_{n=i}^{\infty}s_{n}(W_{n})^{\prime}_{+}(x_{k})$, we will have to find the slope of the linear intervals to the immediate right of $x_{k}$ for all $W_{i}$. We will use $R_{x}$ to represent $W_{i+}^{\prime}(x_{k})$. The first term in the sum will be $R_{x}s_{i}$. Since the derivatives of composed functions will multiply from the chain rule, so the value of the next term is $W^{\prime}_{i+1}(x_{k})=T^{\prime}_{i+1}(W_{i}(x_{k}))R_{x}$. $T_{i+1}$ has two linear segments, giving two slope possibilities to multiply by. The correct one to choose is $-1/(1-a_{i+1})$ because it’s ‘active’ around $x_{k}$ ($x_{k}$ is a peak of $W_{i}$, so $W_{i}(x_{k})>a_{i+1}$ for $x\in(B_{i+1}[k-1],B_{i+1}[k+1])$). This gives $W^{\prime}_{i+1}(x_{k})=-R_{x}\frac{s_{i+1}}{(1-a_{i+1})}$. Note that the second term has the opposite sign as the first.
For all remaining terms, since $x_{k}$ was in $P_{i}$, it is in $V_{j}$ for $j>i$. For $x\in(B_{j+1}[k-1],B_{j+1}[k+1])$, $W_{j}(x)<a_{j+1}$ and the chain rule applies the slope $1/a_{j+1}$. Since this slope is positive, every remaining term continues to have the opposite sign as the first term. Summing up all the terms with the coefficients $s_{i}$, and factoring out $R_{x}$ will yield the desired formula. Note that this same derivation applies to the left sided derivatives as well because the ‘active’ slopes of $T_{i+1}(W_{i})$ are all the same whether a bend in $W_{i}$ is approached from the left or the right. The initial slope constant $L_{x}$ will just be different. ∎
@@
The following lemma shows that to calculate the derivative at of $F(x)$ for any bend point $x$, one needs only to compute the derivative of the finite approximation $f_{i}$ (which excludes $W_{i}$). This will be useful later for proving other results.
@@
From the previous lemma we have $E^{\prime}(x)=0$ whenever the directional derivatives are equal. $F(x)=\sum_{j=0}^{i-1}s_{j}W_{j}(x)+E(x)$. The first $i-1$ terms are differentiable at the points $P_{i}$ since those points lie between the discontinuities in $B_{i-1}$. Therefore $F^{\prime}(x)$ is defined and can be calculated using the finite sum. A visualization of this lemma is provided in Figure 11. ∎
We now prove our main theorem, which shows that there is a way to sum the triangular waveforms $W_{i}$ so that the resulting approximation converges to a continuously differentiable function. The idea of the proof is that much of the formula for $E^{\prime}(x)$ will be shared between two successive generations of peaks. Once they are both valleys, they will be treated the same by the remaining compositions, so the sizes of their remaining discontinuities will need to be proportional.
@@
$F^{\prime}(x)$ is continuous on $[0,1]$ only if the scaling coefficients are selected based on $a_{i}$ according to:
@@
Figure 11: The derivatives of the first few stages of approximation. Notice that each time a constant segment “splits”, the two neighboring segments adjacent to the split monotonically converge back to the original value (marked with a colored point corresponding to the step it’s from)
A.4 Sufficiency for differentiability
This form of the scaling equation derived in the previous section is rather interesting. Since the ratio of two successive scaling terms is $(1-a_{i+1})a_{i+2}$, factors of both $a_{i}$ and $1-a_{i}$ are present in $s_{i}$. This has a few implications, firstly if any $a_{i}$ is 0 or 1, subsequent scales will all be 0, essentially freezing the corresponding neural network at a finite depth. Secondly, having both $a_{i}$ and $1-a_{i}$ will cancel whichever slope multiplier $T_{i}$ contributes to $W_{i}$ at each point $x$, leaving behind the other term, which is less than 1.
If we ensure each $a_{i}$ is bounded away from 0 or 1 by being drawn from the an interval such that $c<a_{i}<1-c$ for some $0.50<c<$ 0.5, then the maximum value of $W^{\prime}_{i}$ is always upper bounded by $(1-c)^{i}$, which is sufficient for uniform convergence. This condition also means that the bends points (and thus the activation regions of the corresponding neural network) will become dense in $[0,1]$, as each region is partitioned (at worst) in a $c$ : $1-c$ ratio.
Lastly, we can show that bounds on $a_{i}$ and our choice of scaling values are sufficient for the existence of $F^{\prime}$ on the bend points, in addition to being necessary for the existence of the derivative on the bend points, our choice of scaling is sufficient when $a_{i}$ are bounded away from 0 or 1.
@@ A.4 Sufficiency for differentiability
on bend points $x$, $F^{\prime}(x)$ exists if we can find $c>0$ such that $c\leq a_{i}\leq 1-c$ for all $i$ and choose all $s_{i}$ according to Equation 4.
@@
We will prove our result by substituting Equation 4 into this formula, and then verifying that the resulting equation is valid. First we would like to rewrite each occurrence of $s$ in terms of $s_{i}$. Equation 4 gives a recurrence relation. Converting it to an non-recursive representation we have:
@@
This equation has a meaningful interpretation that is important to the argument. 1 is the full size of the initial derivative discontinuity at a point in $P_{i}$, and each term on the other side represents proportionally how much the discontinuity is closed for each triangle wave that is added. Every time a wave is added, it subtracts the first term appearing on the right hand side. The following argument shows that each term of the sum on the right accounts for a fraction (equal to $a_{i}$) of the remaining discontinuity, guaranteeing its disappearance in the limit. Inductively we can show:
@@
In words this means that as the first term appearing on the right in Equation 8 is repeatedly subtracted, that term is always equal to $a_{n}$ times the left side. As a base case, we have $(1-a_{i+2})=(1-a_{i+2})$. Assuming the above equation holds for all previous values of $j$
@@ A.5 Error Decay
A.6 Second Derivatives
Here we show that any function represented by one of these networks that is not $y=x^{2}$ does not have a continuous second derivative, as it will not be defined at the bend locations. To show this we will sample a discrete series of $\Delta y/\Delta x$ values from $F^{\prime}(x)$ and show that the limits of these series on the right and left are not the same (unless all $a_{i}=0.5$), which implies that $F^{\prime\prime}(x)$ does not exist (see Figure 12 below). First we will produce the series of $\Delta x$. Let $x$ be the location of a peak of $W_{i}$, and let $l_{n}$ and $r_{n}$ be its immediate neighbors in $B_{i+n}$.
@@ A.6 Second Derivatives
Figure 12: An illustration of attempting to calculate the second derivative. The points in the series approaching $x$ from the right are marked. We rely on the fact that at bend points, the first derivative converges back to the value it had at a finite point in the approximation. $a_{0}\neq 0.5$ and all other parameters are set to $0.5$, which will cause the left and right sets of points to lie on lines with different slopes.
@@
The points $l_{n}$ and $r_{n}$ are all peak locations, Equation 6 gives their derivative values as $f^{\prime}_{i+n}(r_{n})$. In our argument for sufficiency, we reasoned about the sizes of the discontinuities in $f^{\prime}$ at $x$. Since $l_{n}$ and $r_{n}$ always lie on the linear intervals surrounding $x$ as $n\rightarrow\infty$, we can get the value of $f^{\prime}_{i}(x)-f^{\prime}_{i+n}(r_{n})$ using Equation 9 with the initial discontinuity size set to $Rs_{i}$ rather than 1. Focusing on the right hand side we get:
@@
The issue which arises is that the derivation on the left is identical, except for a replacement of $R^{2}$ by $L^{2}$. The only way for these formulas to agree then is for $R^{2}=L^{2}$ which implies $a_{i}=1-a_{i}=0.5$. Since this argument applies at any layer, then all $a$ parameters must be $0.5$ (which approximates $y=x^{2}$). ∎
A.7 Monotonicity and Continuity of Derivatives
Each of the $f^{\prime}_{i}$ are composed of constant value segments, we will show that those values are monotonically decreasing (this can be seen in Figure11). This can extend into the limit to show that $F^{\prime}$ is monotone decreasing and that $F$ is concave.
@@ A.7 Monotonicity and Continuity of Derivatives
We establish two inductive invariants. One is that the y-values in the list $Y_{i}$ remain sorted in descending order. The other is that $Y^{\prime}_{i}[n]\geq f^{\prime}_{i}(x)\geq Y^{\prime}_{i}[n+1]$ for $V_{i}[n]<x<V_{i}[n+1]$, indicating that the constant value segments of $f_{i}$ lie in between the limits in the list $Y_{i}$. Together, these two facts imply that each iteration of the approximation $f_{i}$ is concave. Which then be used to prove that their limit $F$ is also concave.
@@
Lastly we briefly sketch out why $F^{\prime}$ is continuous. It relies on some of our earlier reasoning. Monotonicity of the derivative makes continuity easy to show, because when $x_{1}$ within $\delta$ of $x_{2}$ has $f(x_{1})$ within $\epsilon$ of $f(x_{2})$, so do all intermediate values of $x$. We can establish continuity of $F^{\prime}$ at bend points $x$ easily by using Equation 9. We can pick an $i$ so that the constant value segments of $f^{\prime}_{i}$ are within $\epsilon$ of $F^{\prime}(x)$ and then use the next iteration of bend points (since the constant intervals split, but the new segments near $x_{n}$ converge monotonically towards it) to find $\delta$. In the case of continuity for non bend points $x$, they sit inside a constant-valued interval of $f_{i}$ for each $i$. We can choose $i$ such that $\sum_{n=i}^{\infty}(1-c)^{n}<\epsilon/2$ because this series constrains how far derivative values can move in the limit, and then use the constant interval $x$ is situated in to find $\delta$.
A.8 Activation Region Counting and Zaslavsky’s Theorem
In the coming sections, we present the results of using our 4-neuron-wide construction as an activation function on more complicated datasets. It will be useful to know the impact of this approach on the activation region count when assessing how to best spend a network’s parameter budget. To build up to it, we will provide an informal introduction of Zaslavsky’s theorem (Zaslavsky, 1975). The activation boundaries of ReLU neurons in a single hidden layer network form a hyperplane arrangement. Counting the number of cells, which are convex connected components of space separated by the hyperplanes, is equivalent to determining the number of activation regions produced by the shallow ReLU network. A $D$-dimensional hyperplane arrangement is said to be in general configuration when every set of $D$ hyperplanes intersect in a unique point, every set of $D-1$ hyperplanes intersect in a unique line, etc… General configuration is essentially what you would expect on average from setting the orientations and offsets of the hyperplanes randomly (and would thus be applicable to a single hidden layer ReLU network). Zaslavsky’s theorem states that
Theorem A.10.
When in general configuration, the number of cells ($R\binom{n}{D}$) created by a $D$-dimensional arrangement of $n$ hyperplanes is
$R\binom{n}{D}=\sum_{d=0}^{D}\binom{n}{d}$
To compute this sum, one would go down to the row of Pascal’s triangle corresponding to the number of planes $n$, and sum the first $D$ entries (see Figure 13 for a visual example). It’s a known fact that the sequence of the $d$th entry for each row grows to the $d$th power. So as the number of hyperplanes tends towards infinity, this sum is dominated by the $D$th term, which grows to the power $D$. In other words, the number of linear regions formed by a single hidden layer ReLU network grows exponentially with respect to its input dimension. This property partly explains why neural networks are successful at learning high dimensional functions, even when other classical methods become intractable.
Figure 13: 3 lines dividing 2-dimensional space creates 7 cells. This is calculated by going down to row 3 (for the number of lines) and summing the numbers at indices 0, 1, and 2 (for the dimension). Also note that there is one 2-D plane with 3 lines and 3 points of intersection in this configuration, corresponding (in that order) to the highlighted entries in Pascal’s triangle.
Thanks to the general configuration of the hyperplanes, the number of lower dimensional objects (i.e. points, lines, planes, etc…) are given by the entries in Pascal’s triangle. So a different way to phrase Zaslavsky’s theorem is that the number of cells created by the arrangement is equal to the sum of all these lower dimensional objects, or that each lower dimensional object could be ‘assigned’ a cell. There’s a striking similarity in the intuitions behind the recurrence that generates pascal’s triangle, and the calculation of the numbers of lower dimensional objects created from a hyperplane arrangement. In the case of Pascal’s triangle, if one wished to know how to choose 3 items from 6, one could reason that if item 6 is included, 2 items would be chosen from the remaining 5, and if item 6 is excluded, 3 would be chosen from the remaining 5, and therefore $\binom{6}{3}=\binom{5}{2}+\binom{5}{3}$. Now consider the case of 3 planes partitioning 3-space, and the addition of a 4th plane. The total number of lower dimensional objects would be found by including those that are part of the original 3 planes (1 3-space, 3 planes, 3 lines, 1 point), and adding those that the 4th plane creates (1 plane, 3 lines, 3 points). One could also picture that the new 4th plane ‘cuts’ the existing 3-cells, and to find the number of new cells added, one could examine the projection of the 3 existing planes onto the 4th (shown in Figure 13), and count the 2-cells that appear. Either reasoning leads to $R\binom{n}{D}=R\binom{n-1}{D}+R\binom{n-1}{D-1}$. This similarity between the recurrences means that if the initial conditions can be lined up suitably, Zaslavsky’s theorem can be proven by induction over Pascal’s triangle.
In the case of our architecture, using our 1-dimensional construction as an activation function will generate $n$ generally configured sets of $m$ parallel hyperplanes. To account for this, a slight modification is made to the recurrence. Adding an additional set of $m$ hyperplanes would add $m$ times as many regions, as each of the lower dimensional objects in the existing configuration would project onto each of the $m$ additional hyperplanes. The effect that this has on the formula is that $R\binom{n,m}{D}=R\binom{n-1,m}{D}+m\times R\binom{n-1,m}{D-1}$. Inducting over Pascal’s triangle (with a base case of $R\binom{n,m}{1}=1+nm$) would give the non-recursive formula of $R\binom{n,m}{D}=\sum_{d=0}^{D}m^{d}\binom{n}{d}$. This means that out architecture (with a single hidden layer of depth 5 convex function blocks) would be generating $32^{D}$ times as many regions as an ordinary single layer of ReLU, which is a fantastic trade off once the layers are wide enough (both so that the largest exponent dominates the sum in Zaslavsky’s theorem, and so that only a few neurons would be sacrificed to pay for the new parameters in the activations). Of course, the effect on accuracy will all depend on how well these regions can be utilized.
A.9 Learning Rates